OpenStudy (anonymous):

the three circles are tangent to each other(see fig.),have their centres on the diameter AB of larger circle . if AB=6cm ,then the sum of perimeters of the three small circle are A)pie(22/7)B)4pieC)6piecmD)data insuffient

OpenStudy (anonymous):



OpenStudy (anonymous):

@ganeshie8

OpenStudy (ikram002p):

AB=6 diamater of small circle =6/3=2 perimeter of a small circle=diamater*pi =2pi u have three circles so totall perimeter=3*2*pi=6 pi

ganeshie8 (ganeshie8):